# JWST COMPASS: Insights into the Systematic Noise Properties of NIRSpec/G395H From a Uniform Reanalysis of Seven Transmission Spectra

### v1

@@ Abstract
JWST has already observed near-infrared transmission spectra of over a dozen super-Earths and sub-Neptunes. While some observations have allowed astronomers to characterize sub-Neptunes in unprecedented detail, small feature amplitudes and poorly-understood systematics have led to ambiguous results for others. Using the first seven targets from the COMPASS program, which is currently surveying 12 small planet atmospheres using NIRSpec/G395H, we investigate these timeseries systematics. We implement a model that uses the principle components of the normalized pixel fluxes to account for variations in the shape and position of the spectral trace. We find that observations with a smaller number of groups-per-integration benefit most profoundly from the use of this model, and that systematics are particularly strong between 2.8 and 3.5 m. Despite these systematics, pandexo is a relatively accurate predictor of the precision of the spectra, with real error bars on average 5% larger in NRS1 and 12% larger in NRS2 than predicted. We compute new limits on metallicity and opaque pressure level for each target and compare these to previous results from the COMPASS program. Next, we co-add spectra from multiple targets to reduce the effective noise in the combined spectra in hopes of detecting transmission features in common between the targets, but this exercise does not yield compelling evidence any signals. We find that a handful of additional transits are sufficient to break the degeneracy between metallicity and aerosols for the majority of our targets, pointing towards the possibility of unraveling the mysteries of these worlds with future allocations of JWST time.
1 Introduction
@@
Sub-Neptunes and super-Earths, planets larger than the Earth but smaller than Neptune’s radius of 3.86, are a numerous class of exoplanets that have no counterpart in the Solar System. Since the discovery of this population of small planets by the Kepler mission (Borucki2011; Batalha2013) the origins and atmospheres of these planets have been a mystery. As super-Earths and sub-Neptunes represent potentially habitable environments (Nixon2021; Luger2015) and make enticing targets for biosignature searches (Rauer2011; Seager2021; Tsai2024; Madhusudhan2021; Phillips2021), understanding their histories and present-day environments is a key goal for exoplanet scientists.
@@
The masses and radii of super-Earths and sub-Neptunes permit a range of possible compositions (Seager2007; Fortney2007; Valencia2007). The smallest members of this population, the super-Earths, are likely to have a predominantly rocky interior composition and may or may not possess atmospheres (Ginzburg2016). Throughout this work we use the term “super-Earth” to refer to planets below the radius gap (Fulton2017) at about 1.8, a region of the mass-radius parameter space with a low occurrence of planets which is thought to separate terrestrial worlds from the lower density sub-Neptunes. The radii of sub-Neptunes, defined in this work as any planets smaller than Neptune but above the radius gap, can be explained by several different classes of interior models including gas dwarfs (a rocky core underneath a thick, low mean molecular weight atmosphere, Lopez2014; Rigby2024), true mini-Neptunes (small rocky cores overlain by some combination of water ice, liquid water, and an H2-dominated atmosphere, Rogers2011), steam worlds (a rocky core underneath a high mean molecular weight steam/supercritical H2O atmosphere, Mousis2020; Aguichine2021), or even Hycean worlds (a rocky core overlain by thick hydrosphere with liquid water at the surface and a thin H2-dominated atmosphere, Madhusudhan2021; Madhusudhan2023). While some studies have attempted to deduce the existence of distinct populations of these planet types by careful examination of their mass-radius distribution, these efforts have been hampered by degeneracies between different interior models and uncertainties in models of atmospheric loss (Otegi2020; Rogers2023; Burn2024).
@@
While the NIRSpec instrument has displayed excellent performance and several JWST sub-Neptune observations have allowed for detailed atmospheric characterizations (e.g. Benneke2024; Madhusudhan2023; Schlawin2024), atmospheric signals for all but the nearest planets with clear, low-metallicity, large scale-height atmospheres still sit near the observed noise floor. This has resulted in a number of more ambiguous observations (e.g. Ahrer2025; Cadieux2024, and the COMPASS papers listed in Table 1). Compounding this challenge is the fact that JWST time-series observations frequently display red noise of uncertain origin on timescales of minutes to hours with amplitudes of tens to hundreds of parts-per-million in the white light curves (e.g., Luque2024; Wallack2024; Sarkar2024), degrading the precision of transmission spectra to below what is predicted by tools like Pandeia (Pontoppidan2016) and Pandexo (Batalha2017) (which uses Pandeia as the backend noise simulator). Studies of individual planets observed with NIRSpec have dealt with this noise in a variety of ways, for example by fitting the light curve with a systematics model that incorporates auxiliary timeseries such as the x and y-positions of the trace on the detector as measured during spectral extraction (e.g. Rustamkulov2022; Scarsdale2024) or various engineering parameters associated with the telescope (e.g., Wallack2024; Kirk2024). While these techniques show promise, there is still a need to develop a more complete understanding of detector effects as well as a set of best-practices for mitigating red noise and reaching closer to the photon noise limit. A primary aim of this study is to progress towards that goal with respect to NIRSpec.
@@
In this work we leverage the uniformity of the COMPASS dataset of sub-Neptune and super-Earth transmission spectra to study the characteristics of the NIRSpec detectors and the noise properties of the light curves. All COMPASS targets were observed using NIRSpec/G395H in the Bright Object Time-Series mode with the SUB2048 subarray, F290LP filter, and S1600A1 slit in the NRSRAPID readout pattern. The number of transits for each target was chosen to achieve a transit SNR of 30 ppm at 4 m when binned to a spectral resolution of R=100, according to pre-observation simulations using Pandexo. We carry out uniform reductions of all seven datasets using ExoTiC-JEDI (Alderson2022) holding all reductions procedures and settings identical between all targets, followed by white and spectral light curve fitting using a custom fitting code. We describe both the reduction and the fitting procedure in Section 2. When fitting the light curves we apply a newly-developed systematics model, which is introduced in Section 2 and described in more detail in Section 3. Using this model, we proceed to analyze the instrumental systematics present in NIRSpec/G395H observations and discuss their impacts on transmission spectra.
@@
Following the analysis of the timeseries systematics we detail our methodology for constraining the atmospheric metallicity and aerosol properties of the planets in Section 4. Our method is similar to that used in previous COMPASS papers, with the difference that we use MCMC to sample the likelihood space whereas previous papers used a grid of models in metallicity and opaque pressure level and compared these to the transmission spectra to find the contour along which these models could be ruled out at . In Section 5 we present our lower limits on metallicity and opaque pressure level for each planet. We compare these to the limits obtained in previous papers in order to understand how robust these lower limits are with respect to differences in the reduction and modeling assumptions. In Section 6 we draw comparisons between these seven COMPASS targets and the eleven other planets smaller than 3 observed in transmission with JWST in the near-infrared. We then investigate the potential for detecting transmission features and breaking the metallicity-opaque pressure level degeneracy for the seven COMPASS targets with additional observations, before offering some concluding remarks in Section 7.
2 Observations and Data Reduction
@@
The seven planets included in this study are GJ 357 b (TOI-562.01, Adams2025), TOI-836 b (TOI-836.02, Alderson2024), TOI-836 c (TOI-836.01, Wallack2024), TOI-776 b (TOI-776.02, Alderson2025), TOI-776 c (TOI-776.01, Teske2025), L 98-59 c (TOI-175.01, Scarsdale2024), and L 168-9 b (TOI-134.01, Alam2025). Details of each of these observations can be found in the associated papers, but to summarize, all observations were made with JWST NIRSpec in the G395H mode, which provides spectra between 2.87 and 5.14 m with an average resolution of R=2700. All observations used the NIRSpec Bright Object Time Series mode with the SUB2048 subarray, F290LP filter, and S1600A1 slit in the NRSRAPID readout pattern. NIRSpec/G395H spectra are dispersed across the two NIRSpec detectors, NRS1 and NRS2, resulting in a gap from 3.72 to 3.82 m. Table 1 summarizes the observations for each target.
@@
| GJ 357 b | 1 | 3 | 4401 | No | Adams2025 |
| TOI-836 b | 2 | 3 | 5259 | No | Wallack2024 |
| TOI-836 c | 1 | 3 | 6755 | No | Alderson2024 |
| TOI-776 b | 2 | 7 | 3288 | Transit 2 | Alderson2025 |
| TOI-776 c | 2 | 7 | 3636 | Transit 2 | Teske2025 |
| L 98-59 c | 2 | 4 | 3311 | No | Scarsdale2024 |
| L 168-9 b | 3 | 4 | 3359 | Transit 1 | Alam2025 |
2.1 ExoTiC-JEDI Reduction
We carried out a homogeneous reduction of all observations using the ExoTiC-JEDI package (Alderson2022). By homogeneous reduction, we mean that all parameters of the reduction were held constant between observations. For the first stage of the reduction we use the standard jwst pipeline (version 1.18.1, context map 1364; jwst_pipeline), substituting the custom bias subtraction routine implemented in ExoTiC-JEDI and described in Alderson2024 and adding in the custom group-level destriping routine from ExoTic-JEDI, which uses the column-by-column median value of the out-of-trace pixels to correct for 1/f noise, where the in-trace pixels are identified as all those within standard deviations of the trace center, with the trace center and width defined by a second-order polynomial fit to the column-by-column medians and standard deviations of a Gaussian fit to the PSF. We then conduct ramp fitting using the routine implemented in the jwst pipeline.
@@ 2.1 ExoTiC-JEDI Reduction
We extract the spectrum using the intrapixel aperture extraction routine implemented in ExoTiC-JEDI. The extraction aperture is defined in units of the full-width half-maximum (FWHM) of the trace and is allowed to vary across the detector as the width of the trace increases at longer wavelengths. The extraction aperture for each target, visit, and detector is individually selected by carrying out a trial extraction of the last 500 integrations of the observation for candidate aperture widths ranging from 2 to 10 FWHM in units of 0.2 FWHMs. We chose to carry out this trial extraction on the last 500 integrations in order to avoid selecting in-transit points and to avoid any possible effects of instrument settling at the beginning of the observation. Table 2 contains the extraction apertures in units of FWHM and the corresponding aperture width in pixels at 4 m for each observation.
@@
2.2 White Light curve Analysis
2.2.1 Systematics Model
The origin of the correlated noise seen in NIRSpec time-series observations is uncertain, and it may arise from a combination of both astrophysical and instrumental sources. Previous studies have attempted to mitigate red noise by fitting the light curve with a systematics model that incorporates auxiliary timeseries such as the x and y-positions of the trace on the detector as measured during spectral extraction (e.g. Rustamkulov2022; Alderson2023; Scarsdale2024) or various engineering telemetry timeseries associated with the telescope (e.g., Wallack2024; Kirk2024). We take a similar approach, but rather than using engineering parameters or trace positions we use basis vectors extracted using a principal component analysis (PCA) of the relative pixel flux (RPF) timeseries. principal component analysis is a dimension reduction technique that can, for our purposes, be understood as a method for reducing the full set of RPF vectors to a set of orthogonal vectors that together explain the majority of the variance in the relative pixel fluxes. PCA techniques have been used to reveal detector-level artifacts in Kepler (Luger2018) and Spitzer observations (Deming2015) as well as for NIRISS/SOSS (Radica2025; Coulombe2023) and NIRSpec observations (Luque2024, discussed further below) from JWST.
@@ 2.2.1 Systematics Model
We then compute the principal components of the set of RPF timeseries for each individual visit and detector using the PCA algorithm implemented in SciPy (scipy). PCA is a method for reducing the dimensionality of a dataset, in this case the collection of RPF timeseries, when the individual members of the dataset can be represented as linear combinations of a smaller set of basis vectors plus some additional noise. By applying this technique, we are able to identify a set of principal component timeseries which represent the components of the correlated noise which are shared among the RPF vectors. We construct our systematics model as a linear combination of these principal components, plus a first order polynomial trend in time. The systematics model for the visit is therefore given by
@@
We demonstrate in Section 3 that the principal component vectors correspond to small changes in the morphology of the trace on the detector over time. This systematics model thus extends the practice of fitting shifts in the position of the trace to include more complex, higher-order changes in trace morphology. It has the advantage that the basis vectors are extracted directly from pixel fluxes and is therefore independent of the method of spectral extraction, making the technique pipeline-independent. It can therefore be used with pipelines that carry out optimal extraction such as Eureka! (Bell2022) or aperture extraction such as ExoTiC-JEDI.
@@
Luque2024 constructs a similar systematics model for their NIRSpec/G395H observations of TOI-1685 b’s phase curve. However, their model differs in that it uses the total flux in each pixel rather than the relative pixel flux. By using the relative pixel flux we remove any astrophysical trends while preserving detector effects such a residual 1/f noise, gain variations, and changes in the shape or position of the trace. This potentially explains our model’s success at fitting red noise whereas their model overfit the transit signal and was not found to reduce the red noise. In addition to PCA, Luque2024 also use Independent Component Analysis, or ICA, to decompose the pixel timeseries and obtain a set of basis vectors. ICA is related to PCA, but rather than ensuring that the resulting components are orthogonal, ICA guarantees that the components will be statistically independent. In developing our systematics model we also experimented with ICA, and found that while ICA and PCA produce distinct basis vectors, the best-fit or maximum-likelihood systematics model is indistinguishable between the two techniques.
@@
The phase curve observations for TOI-1685 b use 16 groups per integration, which is larger than the three to seven groups used for the COMPASS observations. As we will see in Section 3, the amplitude of the systematics decreases with the number of groups, so that a 16 group observation may not display significant red noise of the type seen in our observations. This is borne out by the fact that the red noise in the TOI-1685 b observation occurs mainly on timescales of hours, suggesting that it has a different source altogether.
2.2.2 White Light curve Fitting Procedure

### v2

@@ Abstract
JWST has already observed near-infrared transmission spectra of over a dozen super-Earths and sub-Neptunes. While some observations have allowed astronomers to characterize sub-Neptunes in unprecedented detail, small feature amplitudes and poorly-understood systematics have led to ambiguous results for others. Using the first seven targets from the COMPASS program, which will survey 12 small planets using NIRSpec/G395H, we investigate timeseries systematics. We implement a model that uses the principle components of the normalized pixel fluxes to account for variations in the shape and position of the spectral trace. We find that observations with a smaller number of groups-per-integration benefit most profoundly from the use of this model, and that systematics are particularly strong between 2.8 and 3.5 m. Despite these systematics, pandexo is a relatively accurate predictor of the precision of the spectra, with real error bars on average 5% larger in NRS1 and 12% larger in NRS2 than predicted. We compute new limits on metallicity and opaque pressure level for each target and compare these to previous results from the COMPASS program. Next, we co-add spectra from multiple targets to reduce the effective noise in the combined spectra in hopes of detecting transmission features in common between the targets, but this exercise does not yield compelling evidence any signals. We find that a handful of additional transits are sufficient to break the degeneracy between metallicity and aerosols for the majority of our targets, pointing towards the possibility of unraveling the mysteries of these worlds with future allocations of JWST time.
I Introduction
@@
Sub-Neptunes and super-Earths, planets larger than the Earth but smaller than Neptune’s radius of 3.86, are a numerous class of exoplanets that have no counterpart in the Solar System. Since the discovery of this population of small planets by the Kepler mission (Borucki et al., 2011; Batalha et al., 2013) the origins and atmospheres of these planets have been a mystery. As super-Earths and sub-Neptunes represent potentially habitable environments (Nixon and Madhusudhan, 2021; Luger et al., 2015) and make enticing targets for biosignature searches (Rauer et al., 2011; Seager et al., 2021; Tsai et al., 2024; Madhusudhan et al., 2021; Phillips et al., 2021), understanding their histories and present-day environments is a key goal for exoplanet scientists.
@@
The masses and radii of super-Earths and sub-Neptunes permit a range of possible compositions (Seager et al., 2007; Fortney et al., 2007; Valencia et al., 2007). The smallest members of this population, the super-Earths, are likely to have a predominantly rocky interior composition and may or may not possess atmospheres (Ginzburg et al., 2016). Throughout this work we use the term “super-Earth” to refer to planets below the radius gap (Fulton et al., 2017) at about 1.8, a region of the mass-radius parameter space with a low occurrence of planets which is thought to separate terrestrial worlds from the lower density sub-Neptunes. The radii of sub-Neptunes, defined in this work as any planets smaller than Neptune but above the radius gap, can be explained by several different classes of interior models including gas dwarfs (a rocky core underneath a thick, low mean molecular weight atmosphere, Lopez and Fortney 2014; Rigby et al. 2024), true mini-Neptunes (small rocky cores overlain by some combination of water ice, liquid water, and an H2-dominated atmosphere, Rogers et al. 2011), steam worlds (a rocky core underneath a high mean molecular weight steam/supercritical H2O atmosphere, Mousis et al. 2020; Aguichine et al. 2021), or even Hycean worlds (a rocky core overlain by thick hydrosphere with liquid water at the surface and a thin H2-dominated atmosphere, Madhusudhan et al. 2021, 2023). While some studies have attempted to deduce the existence of distinct populations of these planet types by careful examination of their mass-radius distribution, these efforts have been hampered by degeneracies between different interior models and uncertainties in models of atmospheric loss (Otegi et al., 2020; Rogers et al., 2023; Burn et al., 2024).
@@
While the NIRSpec instrument has displayed excellent performance and several JWST sub-Neptune observations have allowed for detailed atmospheric characterizations (e.g. Benneke et al., 2024; Madhusudhan et al., 2023; Schlawin et al., 2024), atmospheric signals for all but the nearest planets with clear, low-metallicity, large scale-height atmospheres still sit near the observed noise floor. This has resulted in a number of more ambiguous observations (e.g. Ahrer et al., 2025; Cadieux et al., 2024, and the COMPASS papers listed in Table 1). Compounding this challenge is the fact that JWST time-series observations frequently display red noise of uncertain origin on timescales of minutes to hours with amplitudes of tens to hundreds of parts-per-million in the white light curves (e.g., Luque et al., 2024; Wallack et al., 2024; Sarkar et al., 2024), degrading the precision of transmission spectra to below what is predicted by tools like Pandeia (Pontoppidan et al., 2016) and Pandexo (Batalha et al., 2017) (which uses Pandeia as the backend noise simulator). Studies of individual planets observed with NIRSpec have dealt with this noise in a variety of ways, for example by fitting the light curve with a systematics model that incorporates auxiliary timeseries such as the x and y-positions of the trace on the detector as measured during spectral extraction (e.g. Rustamkulov et al., 2022; Scarsdale et al., 2024) or various engineering parameters associated with the telescope (e.g., Wallack et al., 2024; Kirk et al., 2024). While these techniques show promise, there is still a need to develop a more complete understanding of detector effects as well as a set of best-practices for mitigating red noise and reaching closer to the photon noise limit. A primary aim of this study is to progress towards that goal with respect to NIRSpec.
@@
In this work we leverage the uniformity of the COMPASS dataset of sub-Neptune and super-Earth transmission spectra to study the characteristics of the NIRSpec detectors and the noise properties of the light curves. All COMPASS targets were observed using NIRSpec/G395H in the Bright Object Time-Series mode with the SUB2048 subarray, F290LP filter, and S1600A1 slit in the NRSRAPID readout pattern. The number of transits for each target was chosen to achieve a transit SNR of 30 ppm at 4 m when binned to a spectral resolution of R=100, according to pre-observation simulations using Pandexo. We carry out uniform reductions of all seven datasets using ExoTiC-JEDI (Alderson et al., 2022) holding all reductions procedures and settings identical between all targets, followed by white and spectral light curve fitting using a custom fitting code. We describe both the reduction and the fitting procedure in Section II. When fitting the light curves we apply a newly-developed systematics model, which is introduced in Section II and described in more detail in Section III. Using this model, we proceed to analyze the instrumental systematics present in NIRSpec/G395H observations and discuss their impacts on transmission spectra.
@@
Following the analysis of the timeseries systematics we detail our methodology for constraining the atmospheric metallicity and aerosol properties of the planets in Section IV. Our method is similar to that used in previous COMPASS papers, with the difference that we use MCMC to sample the likelihood space whereas previous papers used a grid of models in metallicity and opaque pressure level and compared these to the transmission spectra to find the contour along which these models could be ruled out at . In Section V we present our lower limits on metallicity and opaque pressure level for each planet. We compare these to the limits obtained in previous papers in order to understand how robust these lower limits are with respect to differences in the reduction and modeling assumptions. In Section VI we draw comparisons between these seven COMPASS targets and the eleven other planets smaller than 3 observed in transmission with JWST in the near-infrared. We then investigate the potential for detecting transmission features and breaking the metallicity-opaque pressure level degeneracy for the seven COMPASS targets with additional observations, before offering some concluding remarks in Section VII.
II Observations and Data Reduction
@@
The seven planets included in this study are GJ 357 b (TOI-562.01, Adams Redai et al., 2025), TOI-836 b (TOI-836.02, Alderson et al., 2024), TOI-836 c (TOI-836.01, Wallack et al., 2024), TOI-776 b (TOI-776.02, Alderson et al., 2025), TOI-776 c (TOI-776.01, Teske et al., 2025), L 98-59 c (TOI-175.01, Scarsdale et al., 2024), and L 168-9 b (TOI-134.01, Alam et al., 2025). Details of each of these observations can be found in the associated papers, but to summarize, all observations were made with JWST NIRSpec in the G395H mode, which provides spectra between 2.87 and 5.14 m with an average resolution of R=2700. All observations used the NIRSpec Bright Object Time Series mode with the SUB2048 subarray, F290LP filter, and S1600A1 slit in the NRSRAPID readout pattern. NIRSpec/G395H spectra are dispersed across the two NIRSpec detectors, NRS1 and NRS2, resulting in a gap from 3.72 to 3.82 m. Table 1 summarizes the observations for each target.
@@
| GJ 357 b | 1 | 3 | 4401 | No | Adams Redai et al. (2025) |
| TOI-836 b | 2 | 3 | 5259 | No | Wallack et al. (2024) |
| TOI-836 c | 1 | 3 | 6755 | No | Alderson et al. (2024) |
| TOI-776 b | 2 | 7 | 3288 | Transit 2 | Alderson et al. (2025) |
| TOI-776 c | 2 | 7 | 3636 | Transit 2 | Teske et al. (2025) |
| L 98-59 c | 2 | 4 | 3311 | No | Scarsdale et al. (2024) |
| L 168-9 b | 3 | 4 | 3359 | Transit 1 | Alam et al. (2025) |
II.1 ExoTiC-JEDI Reduction
We carried out a homogeneous reduction of all observations using the ExoTiC-JEDI package (Alderson et al., 2022). By homogeneous reduction, we mean that all parameters of the reduction were held constant between observations. For the first stage of the reduction we use the standard jwst pipeline (version 1.18.1, context map 1364; Bushouse et al., 2025), substituting the custom bias subtraction routine implemented in ExoTiC-JEDI and described in Alderson et al. (2024) and adding in the custom group-level destriping routine from ExoTic-JEDI, which uses the column-by-column median value of the out-of-trace pixels to correct for 1/f noise, where the in-trace pixels are identified as all those within standard deviations of the trace center, with the trace center and width defined by a second-order polynomial fit to the column-by-column medians and standard deviations of a Gaussian fit to the PSF. We then conduct ramp fitting using the routine implemented in the jwst pipeline.
@@ II.1 ExoTiC-JEDI Reduction
We extract the spectrum using the intrapixel aperture extraction routine implemented in ExoTiC-JEDI. The aperture width is defined in units of the full-width half-maximum (FWHM) of the trace and is allowed to vary across the detector as the width of the trace increases at longer wavelengths. Both the center and width of the aperture are smoothed by fitting a fourth-order polynomial to the column-by-column trace centers and widths, themselves determined by a Gaussian fit to the counts in each column. The extraction aperture for each target, visit, and detector is individually selected by carrying out a trial extraction of the last 500 integrations of the observation for candidate aperture widths ranging from 2 to 10 FWHM in units of 0.2 FWHMs. We chose to carry out this trial extraction on the last 500 integrations in order to avoid selecting in-transit points and to avoid any possible effects of instrument settling at the beginning of the observation. Each 500-integration trial lightcurve is then smoothed using a Gaussian filter with a width of 30 integrations to remove red noise and the selected aperture is the one that minimizes the remaining scatter, computed as the standard deviation of the smoothed lightcurve. Table 2 contains the extraction apertures in units of FWHM and the corresponding aperture width in pixels at 4 m for each observation.
@@
II.2 White Light curve Analysis
II.2.1 Systematics Model
The origin of the correlated noise seen in NIRSpec time-series observations is uncertain, and it may arise from a combination of both astrophysical and instrumental sources. Previous studies have attempted to mitigate red noise by fitting the light curve with a systematics model that incorporates auxiliary timeseries such as the x and y-positions of the trace on the detector as measured during spectral extraction (e.g. Rustamkulov et al., 2022; Alderson et al., 2023; Scarsdale et al., 2024) or various engineering telemetry timeseries associated with the telescope (e.g., Wallack et al., 2024; Kirk et al., 2024). We take a similar approach, but rather than using engineering parameters or trace positions we use basis vectors extracted using a principal component analysis (PCA) of the relative pixel flux (RPF) timeseries. Principal component analysis is a dimension reduction technique that can, for our purposes, be understood as a method for reducing the full set of RPF vectors to a set of orthogonal vectors that together explain the majority of the variance in the relative pixel fluxes. PCA techniques have been used to reveal detector-level artifacts in Kepler (Luger et al., 2018) and Spitzer observations (Deming et al., 2015) as well as for NIRISS/SOSS (Radica et al., 2025; Coulombe et al., 2023) and NIRSpec observations (Luque et al., 2024, discussed further below) from JWST.
@@ II.2.1 Systematics Model
We then compute the principal components of the set of RPF timeseries for each individual visit and detector using the PCA algorithm implemented in SciPy (Virtanen et al., 2020). PCA is a method for reducing the dimensionality of a dataset, in this case the collection of RPF timeseries, when the individual members of the dataset can be represented as linear combinations of a smaller set of basis vectors plus some additional noise. By applying this technique, we are able to identify a set of principal component timeseries which represent the components of the correlated noise which are shared among the RPF vectors. We construct our systematics model as a linear combination of these principal components, plus a first order polynomial trend in time. The systematics model for the visit is therefore given by
@@
We demonstrate in Section III that the principal component vectors correspond to small changes in the morphology of the trace on the detector over time. This systematics model thus extends the practice of fitting shifts in the position of the trace to include more complex, higher-order changes in trace morphology. It has the advantage that the basis vectors are extracted directly from pixel fluxes and is therefore independent of the method of spectral extraction, making the technique pipeline-independent. It can therefore be used with pipelines that carry out optimal extraction such as Eureka! (Bell et al., 2022) or aperture extraction such as ExoTiC-JEDI.
@@
Luque et al. (2024) constructs a similar systematics model for their NIRSpec/G395H observations of TOI-1685 b’s phase curve. However, their model differs in that it uses the total flux in each pixel rather than the relative pixel flux. By using the relative pixel flux we remove any astrophysical trends while preserving detector effects such a residual 1/f noise, gain variations, and changes in the shape or position of the trace. This potentially explains our model’s success at fitting red noise whereas their model overfit the transit signal and was not found to reduce the red noise. In addition to PCA, Luque et al. (2024) also use Independent Component Analysis, or ICA, to decompose the pixel timeseries and obtain a set of basis vectors. ICA is related to PCA, but rather than ensuring that the resulting components are orthogonal, ICA guarantees that the components will be statistically independent. In developing our systematics model we also experimented with ICA, and found that while ICA and PCA produce distinct basis vectors, the best-fit or maximum-likelihood systematics model is indistinguishable between the two techniques.
@@
The phase curve observations for TOI-1685 b use 16 groups per integration, which is larger than the three to seven groups used for the COMPASS observations. As we will see in Section III, the amplitude of the systematics decreases with the number of groups, so that a 16 group observation may not display significant red noise of the type seen in our observations. This is borne out by the fact that the red noise in the TOI-1685 b observation occurs mainly on timescales of hours, suggesting that it has a different source altogether.
II.2.2 White Light curve Fitting Procedure